TJHSST Senior Research Project Project Breakdown 2006-2007

Alex Katkova

September 12, 2006

1 Languages

Since this is a modeling-type program, and I’m moderately comfortable working with Java, I’m going to use MASON to create it. I’m not sure yet how to use MASON, so part of the project will be to find out how.

2 1st Quarter

By the end of first quarter I’m hoping to finish the basic outline of the program and get it to display the two floors of TJ, including the walls. If I finish with that, the program could display a dot for each student and teacher at TJ.

3 2nd Quarter

During this quarter, I plan to develop the probablities that determine the locations of the students and teachers at given times of day. I also have to make sure that the students gradually disappear off campus at the end of the school day.

4 3rd Quarter

If the program works as planned then I would add in a user interface that would allow for the altering of variables such as the length of the school day and the length of classes. I would also base the location of the students on their class.